### v1

# Localized Band Gap Resonances in Defect-embedded Monoatomic Chains

Abstract
Historically, structural defects in crystalline material lattice structures have been viewed as a liability, causing the material behavior to deviate from the expected norm. However, the gradual improvement in understanding the role of defects has overturned this view, encouraging positive outlooks, e.g., deliberately engineering defects in structures, that have laid the foundation for modern applications, e.g., silicon chip manufacturing, battery engineering, fuel cell engineering. In recent years, acoustic metamaterials (AMMs), which have emerged as a platform for customizing acoustic behavior, have also embraced this shift in viewpoint, enabling design propositions involving precisely embedded defects in the AMM architecture that may be relevant for applications, e.g., energy harvesting, waveguiding, flow control. The fundamental utility of defects stems from their ability to robustly localize energy at a prescribed location with desirable spatio-temporal characteristics. However, the bulk of existing defect-based AMM studies primarily present a qualitative numerical analysis, prompting the need for an analytical approach quantifying the defect eigenmode characteristics, e.g., frequency, mode shape, resonance velocity amplitude envelope to precisely engineer defects in AMMs. This article presents a modified analytical method to determine the eigenmode characteristics of a defect-embedded grounded monoatomic chain (DMoA) and demonstrates the benefits of defect modes through two DMoA-based AMMs that can accommodate – (i) a virtual ground, and (ii) achieve customized acoustic interaction and absorption. The proposed strategy can be readily extended to more elaborate AMMs and augments the design space for defect-based AMMs.
Keywords
Metamaterials; bandgap resonances; defect modes; energy localization
I Introduction
Historically, the presence of defects or impurities in the crystalline lattice structures of natural materials has been considered an undesirable characteristic that can adversely impact the material properties. Consequently, the materials research community has developed numerous experimental techniques, e.g., X-ray diffraction[1], ultrasound non-destructive evaluation[2, 3], in the past century, to precisely identify the location, and study the impact of these structural (e.g., point[4], line[5]) defects within the material architecture on the meso- and macro-scale dynamics. The systematic progression in the study of defects has elucidated their critical role in determining material performance and gradually shifted the long-standing negative perception associated with them to a point where defects are now viewed as control parameters in materials design. In recent years, the idea of precisely engineering defects into a solid microstructure has gained tremendous traction forming the cornerstone of multiple modern fields of research, e.g., silicon-based chip manufacturing[6], battery science[7], solid-oxide fuel cells[8].
Simultaneously, in the field of mechanics and materials, the advent of acoustic metamaterials (AMMs), primarily featuring periodic, cleverly architected mesoscale building blocks (i.e., unit cells) have offered an invaluable capability to mimic the ordered atomic architectures of natural materials at a more accessible scale and tailor the macroscale material behavior. AMMs have featured diverse unit cell designs, leveraging active elements[9, 10], geometric instabilities[11, 12], multistability[13, 14], and kinematic amplification mechanisms[15, 16], eliciting peculiar material behavior, e.g., auxeticity[17], non-reciprocity[10], directional wave-guiding[18]. Though AMM research is relatively recent, the community has shared identical views with the larger materials research community about the presence of structural or material defects[19]. Similar to the prevalent perception of defects, this view has continuously evolved and, in recent years, has led to the exploration of material architectures with precisely embedded defects to robustly localize energy at desirable locations and spatio-temporal scales. Multiple strategies[20, 21, 22, 23, 24] exist in the literature, to introduce defect modes in a variety of material architectures, e.g., Kaina et al.[20] experimentally demonstrated a waveguide allowing wave propagation at reduced group velocities by altering the lattice spacing of the periodic sub-wavelength resonant structures, Li et al.[21] induced a defect mode in a soft bilayer by triggering a wrinkling instability at a prescribed location via strain-engineering, Xu-Feng et al.[22] experimentally demonstrated high energy localization and efficient energy harvesting by inducing geometric defects in a phononic crystal (PnC) beam. These proposals primarily present a qualitative analysis of the defect resonance dynamics via numerical simulations, parameter studies, supercell analysis, and experiments.
Consequently, subsequent studies have explored analytical methods[25, 26, 27, 28, 29, 30, 31, 32, 33, 34] to characterize the defect mode frequency in terms of the defect parameters, e.g., geometric parameters, material properties. Though these approaches have aided the development of more informed material design strategies, the frequency represents only a subset of the relevant dynamic characteristics of defect modes. Additional modal characteristics, e.g., mode shape, localized velocity (displacement) at the defect location, can also play a significant role in determining the effectiveness of defect-based AMMs and augment their functionality for wave-guiding[35], energy harvesting[36, 37], flow control[38] applications, motivating the need for a more comprehensive characterization of defect modes. Addressing this demand, we present a modified analytical approach to quantify three critical characteristics of defect modes, i.e., frequency, mode shape, and resonance velocity amplitude envelope, in a defect-embedded grounded monoatomic chain.
The remainder of the article is organized as follows – Sec. II provides a comprehensive analysis of the acoustic characteristics of the emergent defect resonances in a defect-embedded grounded monoatomic chain. Sec. III presents two defect-embedded AMMs described in Sec. II, demonstrating the benefits of defect modes. Finally, Sec. IV presents a summary and the key takeaways from the article.
II Model Description
Figure 1: Defect-embedded grounded monoatomic chain (DMoA). (a) Mass-spring model for a defect-free, uniform grounded monoatomic chain (MoA), dispersion curves with pass bands (PBs), band gaps (BGs), and eigenfrequencies of finite ($\mathrm{N}=11$) edge and bulk DMoAs. Mass-spring models and frequency response functions (FRFs) of a finite (b) Edge DMoA, and (c) Bulk DMoA.
Consider a periodic mass-spring model of a 1D grounded monoatomic chain (MoA) as shown in Fig. 1a with mass, $m$, interaction stiffness, $k$ and onsite grounding stiffness, $k_{\mathrm{g}}$. The non-dimensional equation of motion for the $j^{\mathrm{th}}$ mass in an infinite MoA is given by,
$v_{j,\bar{t}\bar{t}}+\left(2v_{j}-v_{j+1}-v_{j-1}\right)+k_{\mathrm{g,r}}v_{j}=0,$ (1)
where, $v_{j}$ is the axial displacement of the $j^{\mathrm{th}}$ mass, $\bar{t}=\omega_{0}t$ is the non-dimensional time, $\omega_{0}=\sqrt{k/m}$ is the reference temporal frequency and $k_{\mathrm{g,r}}=k_{\mathrm{g}}/k$ is the grounding stiffness ratio. Consequently, the dispersion relation for the MoA, assuming a plane wave ansatz of the form $v_{j}=\tilde{v}_{j}e^{i(\kappa x_{j}-\Omega\bar{t})}\ \forall\ j=1,2,\cdots,\mathrm{N}\ \mathrm{and}\ x_{j}=j-1$, is calculated as[39, 40]:
$\Omega(\kappa)=\sqrt{k_{\mathrm{g,r}}+4\sin^{2}\left(\frac{\kappa}{2}\right)},$ (2)
where, $\Omega$ is the non-dimensional temporal frequency, and $\kappa=\kappa_{R}+i\kappa_{I}\ \forall\ \kappa_{R},\kappa_{I}\in\mathbb{R}$, is the non-dimensional complex wavenumber. The dispersion curves (Eq. (2), $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$) and the natural eigenfrequencies, $\Omega_{j}$ of a finite MoA with $\mathrm{N}$ degrees of freedom and free boundaries are plotted in Fig. 1a, highlighting the band gap (BG, $\kappa_{R}=0\ \mathrm{or}\ \pi,\kappa_{I}\neq 0$) and the pass band (PB, $\kappa_{R}\neq 0,\kappa_{I}=0$) frequency ranges. The grounding elastic springs are the primary source of the bounded BG between $\Omega=0$ and a finite $\Omega=\sqrt{k_{\mathrm{g,r}}}$. The grounding springs resist the deformation of masses with respect to a fixed ground and participate in all the eigenmodes describing the MoA dynamics. Consequently, the PB range and specifically the lower edge of the PB, $\Omega(\kappa=0)=\sqrt{k_{\mathrm{g,r}}}$ is directly linked to the elasticity of these grounding elements, opening up a BG below this modal frequency.
When the periodicity of a finite MoA is altered, by introducing a change in one or more material properties (defect), i.e., $m_{\mathrm{r}}=m_{\mathrm{d}}/m$, $k_{\mathrm{r}}=k_{\mathrm{d}}/k$, or $k_{\mathrm{gd,r}}=k_{\mathrm{gd}}/k$ at a specified location within the MoA, the eigenfrequency (or multiple eigenfrequencies, see Sec. A2) associated with the lower (upper) edge of the PB migrates to the lower (upper) BG, creating a BG resonance at $\Omega_{1}=\Omega_{\mathrm{d}}<\sqrt{k_{\mathrm{g,r}}}$ ($\Omega_{\mathrm{N}}=\Omega_{\mathrm{d}}>\sqrt{4+k_{\mathrm{g,r}}}$). Note that a BG resonance can exist even in an ungrounded mass-spring chain[30, 28, 41] however, a grounded chain ($k_{\mathrm{g,r}}\neq 0$) provides an opportunity to engineer a BG resonance in the lower BG, $\Omega_{\mathrm{d}}\in[0,\sqrt{k_{\mathrm{g,r}}})$, that manifests as the first eigenmode of the altered MoA. The existence of a BG resonance as the principal, low-frequency eigenmode may be beneficial in ensuring their effective participation in the system dynamics, facilitating DMoA-based AMMs for practical applications that can involve complex operational requirements, e.g., response to low-frequency broadband external forces[41], amplitude-phase requirements[38]. We refer to the altered MoA as a defect-embedded grounded monoatomic chain (DMoA) and the resultant BG resonance as a defect mode for brevity.
The DMoA can be explored in two fundamentally distinct configurations – (i) an edge DMoA where material properties at one of the boundaries of the finite MoA are altered and, (ii) a bulk DMoA[32] where material properties in the bulk deviate from the periodic properties of the MoA. The natural frequencies of these DMoAs are obtained by solving the matrix eigenvalue problem, $\mathbf{K}_{\mathrm{d}}\hat{\mathbf{V}}=\mathbf{0}$ (see Sec. A1), with $m_{\mathrm{r}}=k_{\mathrm{r}}=1$ and $k_{\mathrm{gd,r}}=0$ and, overlayed onto the dispersion plot in Fig. 1a, indicating the presence of a defect mode (BG resonance) in both cases albeit, with distinct frequencies. In addition, the frequency response functions (FRFs) of the DMoAs, evaluated by applying an external force and monitoring the displacement at the defect mass are plotted in Fig. 1b,c. In either case, the DMoA exhibits relatively higher FRF amplitudes at the defect frequency, $\Omega_{\mathrm{d}}$ compared to other eigenfrequencies, indicating energy localization at the defect mass owing to the manifestation of an exponentially decaying eigenmode shape away from the defect location (see insets of Fig. 1b,c). The remainder of the article uses – $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$, $m_{\mathrm{r}}=1$, $k_{\mathrm{r}}=1$ and $k_{\mathrm{gd,r}}=0$ unless explicitly stated otherwise.
The defect frequency, $\Omega_{\mathrm{d}}$, and the exponential decay rate, $\kappa_{I}$ comprehensively describe the contribution of the material defect in altering the base MoA dynamics. In addition, although rarely considered in prior studies, we quantify and describe a third critical parameter – the velocity amplitude envelope at resonance, $\dot{\mathrm{V}}_{\mathrm{E}}$ that can promote defect-based AMMs for novel applications, e.g., energy harvesting[36, 37], flow control via fluid-metamaterial interactions[38, 42], where response amplitude of the AMM can play a major role in determining operational efficiency. In the following sections, we elaborate on a modified analytical approach to quantify these critical characteristics of the defect dynamics of the DMoA – $\{\Omega_{\mathrm{d}},\kappa_{I},\dot{\mathrm{V}}_{\mathrm{E}}\}$. The proposed approach can be readily extended to general boundary conditions[28, 30] and mass-spring AMM models comprising unit cells with more degrees of freedom, e.g., defect-embedded diatomic[26, 28] and polyatomic[29] chains.
II.1 Defect eigenfrequency and eigenmode shape
Figure 2: Iso-curves of the defect eigenmode characteristics. (Center) 3D plot of the iso-surfaces of defect frequency, $\Omega_{\mathrm{d}}=\{0.4,0.6,0.8,1.0,1.2\}$ as a function of material properties, $m_{\mathrm{r}}$ (log-scale) and $k_{\mathrm{gd,r}}$ and, the size of the base MoA, $\mathrm{N}$. (a) Contour depicting a 2D slice of the 3D central figure, taken at $m_{\mathrm{r}}=1$, (b) Contour depicting a 2D slice of the 3D central figure, taken at $k_{\mathrm{gd,r}}=0$, (c-i) Contour depicting a 2D slice of the 3D central figure, taken at $\mathrm{N}=21$, (c-ii) Contour plotting the analytical solution for defect frequency, $\Omega_{\mathrm{d}}$ as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$, (c-iii) Numerical and (c-iv) analytical contours of the defect eigen mode shape, i.e., decay rate, $\kappa_{I}$ as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. The $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers indicate DMoA configurations investigated in Fig. 3.
Consider the edge DMoA model illustrated in Fig. 1b. The edge defect resembles a local resonator appended to the boundary of a finite, homogeneous elastic foundation (base MoA). The presence of the edge defect creates a localized defect mode in the altered material configuration that can be modeled as a reduced order edge DMoA illustrated in Fig. 1b. The BG eigenfrequencies of a finite edge DMoA are numerically calculated and plotted as iso-defect frequency surfaces, i.e., $\Omega_{\mathrm{d}}=\mathrm{const.}$, in the central 3D-plot in Fig. 2 for varying degrees of freedom, $\mathrm{N}$, and material properties at the edge defect, i.e., $k_{\mathrm{gd,r}}$ and $m_{\mathrm{r}}$, to understand the qualitative dependence of the defect frequency, $\Omega_{\mathrm{d}}$ on these parameters. Fig. 2a presents a 2D slice of the central plot taken at $m_{\mathrm{r}}=1$. Consistent with the qualitative relation between frequency and stiffness, the magnitudes of the iso-defect frequency curves on the contour plot indicate that, $\Omega_{\mathrm{d}}$ increases as $k_{\mathrm{gd,r}}$ increases from $0\longrightarrow k_{\mathrm{g,r}}$. Similarly, in Fig. 2b (2D slice taken at $k_{\mathrm{gd,r}}=0$) the magnitude of iso-defect frequency curves decrease with increasing $m_{\mathrm{r}}$ ($>1$), as anticipated qualitatively. In addition, we notice that the iso-defect frequency curves in both Fig. 2a,b tend to asymptote to a constant value of $k_{\mathrm{gd,r}}$ and $m_{\mathrm{r}}$, respectively, at large $\mathrm{N}$. This observation provides valuable insight for our analytical model, where we can arbitrarily choose a large DMoA size ($\mathrm{N}$) to obtain an accurate prediction for $\Omega_{\mathrm{d}}$, independent of the actual size of the finite DMoA.
Finally, Fig. 2(c-i,iii) plots the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$, calculated from matrix eigen-analysis of a finite DMoA ($\mathrm{N}=21$) as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. A closer look at the matrix eigenvalue problem reveals a tridiagonal matrix form for the dynamic stiffness, $\mathbf{K}_{\mathrm{d}}$ (Sec. A1) that is consistent with the general form of a perturbed tridiagonal Toeplitz matrix[43, 28]. The characteristic equation for $\mathbf{K}_{\mathrm{d}}$ can be written in the implicit form:
$f(\Omega)=g(\Omega),$ (3)
where, without loss of generality for $\mathrm{N}=2n+1$ and $n\in\mathbb{N}$,
$f(\Omega)=\frac{\sin\left([n+1]\left[\cos^{-1}\left(\frac{d^{2}}{2a^{2}}-1\right)\right]\right)}{\sin\left(n\left[\cos^{-1}\left(\frac{d^{2}}{2a^{2}}-1\right)\right]\right)},$ (4)
$g(\Omega)=-\frac{\alpha\beta d+(\alpha+\beta)a^{2}}{(d+\alpha+\beta)a^{2}},$
and,
$d=2+k_{\mathrm{g,r}}-\Omega^{2},a=-1,\beta=-1,$ (5)
$\alpha=(k_{\mathrm{gd,r}}-k_{\mathrm{g,r}})-1+(1-m_{\mathrm{r}})\Omega^{2},$
represent the matrix entries of $\mathbf{K}_{\mathrm{d}}$. Eq. (3) can be solved numerically, assuming a large value of $n$ (i.e, large $\mathrm{N}$) to estimate $\{\Omega_{\mathrm{d}},\kappa_{I}\}$. Fig. 2(c-ii,iv) plot the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ contours, respectively, obtained using the analytical approach, showing an excellent correlation with their numerical counterparts in Fig. 2(c-i,iii). Note that the decay rate, $\kappa_{I}$ only depends on the base MoA properties $k_{\mathrm{g,r}}$ and defect frequency, $\Omega_{\mathrm{d}}$. Therefore, each $\Omega_{\mathrm{d}}$ is associated with a unique $\kappa_{I}$ leading to identical shapes for iso-defect frequency and iso-decay rate curves in Fig. 2c.
Alternately, for the bulk DMoA (Fig.1c), we utilize the U-transformation method[44] proposed by Cai et al.[32] to arrive at an implicit equation identical to Eq. (3) that can be solved numerically to estimate the defect frequency and mode shape. Though an explicit analytical form of $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ remains elusive, the proposed solution approach for both DMoAs require a numerical calculation only at the terminal step to solve a well-defined implicit analytical equation (Eq. (3)). Therefore, the analytical approach executed with a numerical algorithm[45, 46] of our choice allows a systematic derivation of DMoA material parameters that produce the desired spatio-temporal defect characteristics.
II.2 Velocity amplitude envelope at defect mode resonance
Figure 3: Defect Mode shape and response of edge defect at resonance. (a) Normalized velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}(m_{\mathrm{r}},k_{\mathrm{gd,r}})/\dot{\mathrm{V}}_{\mathrm{E}}(1,0)$ contour plot with iso-velocity amplitude envelope curves and the iso-defect frequency curve at $\Omega_{\mathrm{d}}=1$ from Fig. 2(c-i).(b) Defect mode shapes for the $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ pairs indicated by the respective markers, lying on the $\Omega_{\mathrm{d}}=1$ and $\kappa_{I}=0.962$ in Fig. 2c, normalized by the edge defect displacement, $v_{1}(\bar{t})\equiv v_{\mathrm{d}}(\bar{t})$ , (c-i) Velocity response of the edge defect, $\dot{v}_{\mathrm{d}}(\bar{t})$ to a resonant force, $\mathrm{F}(\bar{t})=\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t})$, (c-ii) Magnified view of the numerical and analytical velocity response (Eq. (6)) for $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}=\{1,0.382\}$.
A closer look at the iso-defect frequency and iso-decay rate curves in Fig. 2c indicates the presence of multiple $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ pairs that can manifest the same $\{\Omega_{\mathrm{d}},\kappa_{I}\}$. Therefore, we formulate and study an additional quantity associated with resonance dynamics at the defect location, i.e., the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ (see Sec. A3) to distinguish the dynamic behavior of edge DMoAs with distinct material properties, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. Absolute velocity (displacement) amplitudes at resonance have rarely been featured in fundamental research studies of AMMs. However, their emerging relevance in determining AMM efficiency for applications, e.g., energy harvesting[36, 37], flow control via fluid-metamaterial interactions[38, 42] motivates their consideration as a critical behavioral parameter describing the AMM dynamics. Hence, a prescribed $\dot{\mathrm{V}}_{\mathrm{E}}$ is considered in this section as the third and final parameter to complete the edge DMoA design.
Fig. 3a plots the normalized velocity amplitude envelope (Eq. (A12)) as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. The iso-defect frequency curve at $\Omega_{\mathrm{d}}=1$ intersects the iso-velocity amplitude envelope curves at a single point in the given range of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. Consequently, optimizing for a desired $\{\Omega_{\mathrm{d}},\kappa_{I},\dot{\mathrm{V}}_{\mathrm{E}}\}$ is a well-posed problem that will produce a unique solution, e.g., the $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers indicating $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ values corresponding to $\Omega_{\mathrm{d}}=1$, $\kappa_{I}=0.962$, and $\dot{\mathrm{V}}_{\mathrm{E}}/\dot{\mathrm{V}}_{\mathrm{E}}(1,0)=\{0.96,0.61,0.5,0.39\}$, respectively.
As anticipated, the mode shapes associated with DMoAs described by the $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers, plotted in Fig. 3b manifest an identical decay rate, $\kappa_{I}=0.962$. Each DMoA is now subject to a sinusoidal external force, $\mathrm{F}(\bar{t})=\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t})$ at the edge defect mass, and Fig. 3(c-i) plots the corresponding numerical and analytical velocity response,
$\dot{v}_{\mathrm{d}}(\bar{t})=\dot{\mathrm{V}}_{\mathrm{E}}\bar{t}\sin\left(\Omega_{\mathrm{d}}\bar{t}\right),$ (6)
for a total simulation time is $\bar{t}_{\mathrm{sim}}=1000$.
Consistent with a typical response at resonance, all DMoAs exhibit an oscillatory velocity response at the prescribed defect frequency, $\Omega_{\mathrm{d}}=1$ with a linearly increasing velocity amplitude in time. However, each DMoA manifests a unique rate of increase in velocity amplitude that we designate as the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$. $\dot{\mathrm{V}}_{\mathrm{E}}$ can be analytically estimated in terms of the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ or $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$, and the input force amplitude, $\mathrm{F}_{0}$ (see Sec.A3). Fig. 3(c-ii) contrasts the analytical resonance solution in Eq. (6) and the velocity response recorded in the numerical simulations ($\square$ - marker, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}=\{1,0.382\}$), showing an excellent correlation between the two. Therefore, the proposed analytical approach offers a broad set of design parameters to construct a customized DMoA based on operational requirements.
II.3 Multi-defect embedded grounded monoatomic chain
Figure 4: Multi-defect embedded MoA (Multi-DMoA). (a) Absolute Percentage error contours for a Multi-DMoA ($\mathrm{N}=21$) with two defects at $x_{\mathrm{d1}},x_{\mathrm{d2}}\in[0,\mathrm{N}-1]$ and $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$. (b) Edge and bulk defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$, respectively, of a Multi-DMoA with an edge defect at $x_{\mathrm{d1}}=0$ and a bulk defect at $x_{\mathrm{d2}}\in[1,\mathrm{N}-2]$. FRFs of a Multi-DMoA denoted by
$\times$
-marker in (a,b) when the (c-i) edge and (c-ii) bulk defect masses are excited and, their displacement response monitored, (c-iii) eigenmode shapes of the Multi-DMoA.
In Sec. II.1 and Sec. II.2 we have established the accuracy of the proposed analytical approach in determining the dynamic characteristics of a DMoA with a single defect. But, what are the limitations of the analytical approach? Can the approach be used to characterize DMoAs with multiple defects? To address these questions and investigate the robustness of the solution procedure, in this section, we contrast the numerical and analytical dynamic characteristics of a multi-defect embedded grounded monoatomic chain (Multi-DMoA).
Fig. 4a plots the absolute percentage error between the analytical and numerical values of the two defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}(>\Omega_{\mathrm{d1}})$ generated due to the presence of two defects at locations, $x_{\mathrm{d1}}$ and $x_{\mathrm{d2}}$ in a Multi-DMoA ($\mathrm{N}=21$). The diagonal direction in both these plots remains unpopulated as we assume distinct locations for the embedded material defects, i.e., $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$. The sea of dark blue, low error values ($0.0005\%-2\%$) in both plots indicate the validity of the analytical approach derived in Sec. II.1 for a large section of the Multi-DMoA configurations provided, the two defects in the Multi-DMoA remain relatively isolated from each other. Alternately, as the defect locations become proximal, i.e., within 3 masses, marked by the off-diagonal lines, $x_{\mathrm{d2}}=x_{\mathrm{d1}}\pm 3$, the Multi-DMoA configuration deviates from the isolated (single) defect assumption giving rise to a region of high error ($>2\%$, max. error $\sim 33\%$). However, note that the expanse of the high error region is dependent on the decay rate, $\kappa_{I}$ associated with bulk defect mode, i.e., as the $\kappa_{I}$ grows larger, the high error region shrinks (Sec. A4). Therefore, a modified analytical approach is required to investigate the design space of proximal defects. This lies outside the scope of the current study and isn’t pursued further. Although, prior efforts have studied the dispersion characteristics of phononic crystals with multiple proximal defects via super-cell analysis[26], that can provide good qualitative insights into a potential solution approach.
Another important observation to highlight in the $\Omega_{\mathrm{d1}}$ error plots in Fig. 4a, is the linear light blue regions - $x_{\mathrm{d1}}=1,\mathrm{N}-2$ and $x_{\mathrm{d2}}=1,\mathrm{N}-2$. These lines correspond to Multi-DMoA configurations where one of the penultimate masses is a defect. Technically this defect can be categorized as a bulk defect, however, the proximity to the finite boundary of the Multi-DMoA has a significant impact on the eigenmode characteristics, leading to a slightly higher error ($\sim 3.9\%$) compared to the Multi-DMoA configurations in the dark blue regions. The impact of boundary effects is less prevalent in cases where the defect frequencies lie in the $2^{\mathrm{nd}}$ BG of the MoA (see Sec. A4). We can attribute this phenomenon to the real part of the wavenumber, $\kappa_{R}$ of these defect modes. The defect mode in the $1^{\mathrm{st}}$ BG has $\kappa_{R}=0$ that requires the boundary adjacent to the defect location to vibrate in-phase with the defect at resonance, leading to a rigid body-like motion of the boundary and defect masses. Alternately, the $2^{\mathrm{nd}}$ BG requires $\kappa_{R}=\pi$, thus, allowing the boundary mass to vibrate out of phase with the defect leading to relatively decoupled, anti-symmetric dynamics between the sites rendering the actual defect frequency for the latter case closer to the analytical estimate than the former.
Now, consider the configurations described by the y-axis of the contour plots in Fig. 4a that represents a Multi-DMoA with an edge ($x_{\mathrm{d1}}=0$) and a bulk ($x_{\mathrm{d2}}\in[1,\mathrm{N}-2]$) defect. Fig. 4b plots the numerical edge and bulk defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$, respectively, for different locations of the bulk defect, $x_{\mathrm{d2}}$ in the Multi-DMoA. Consistent with the error plots, both $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$ asymptote to the respective edge-bulk analytical estimates as the bulk defect migrates farther from the edge defect. To further explore interesting dynamics, we consider the Multi-DMoA described by the $\times$ -marker in Fig. 4a,b. Fig. 4c-i (Fig. 4c-ii) plots the FRF when the edge (bulk) defect is subject to a harmonic excitation and its displacement monitored. The relatively low amplitude and monotonic FRF curves when exciting the edge (bulk) defect mass at the other defect frequency, $\Omega_{\mathrm{d2}}$ ($\Omega_{\mathrm{d1}}$) presents an interesting insight in Fig. 4c-i (Fig. 4c-ii). The edge and bulk defect mode shapes plotted in Fig. 4c-iii explain this peculiar dynamical behavior of the Multi-DMoA. In the bulk defect mode, $x_{\mathrm{d2}}=13$ constitutes an anti-node, while $x_{\mathrm{d1}}=0$ constitutes a node. Consequently, exciting the edge defect mass ($x_{\mathrm{d1}}=0$) at $\Omega_{\mathrm{d2}}$ fails to excite the bulk defect mode, and exciting the bulk defect mass ($x_{\mathrm{d2}}=13$) at $\Omega_{\mathrm{d1}}$ fails to excite the edge defect mode, leading to a relatively muted dynamical response at the respective excitation locations in both scenarios. Therefore, the analytical approach provides a robust estimate of the defect mode frequencies, mode shapes, and the velocity amplitude envelope even in the presence of multiple defects in the DMoA, provided the additional defects are nodal points in the eigenmode associated with the considered defect.
III Defect-embedded Acoustic Metamaterials
In this section, we demonstrate how tailored edge DMoA-based AMMs can – (i) overcome the absence of an inertial frame of reference by accommodating an in-house virtual ground within the AMM structure, and (ii) manifest a customized amplitude and narrow-band frequency in response to broadband acoustic fields, e.g., acoustic waves generated due to vehicular traffic, dynamic pressure field of a fluid which may be relevant to applications, e.g., noise/vibration mitigation, energy harvesting, flow control.
Note, we include nominal viscous damping, $c_{\mathrm{r}}=c/\sqrt{km}=0.01$ in numerical simulations presented in this section, to realize bounded acoustic responses and test the robustness of our analytical model.
III.1 Acoustic Metamaterial with a virtual ground
Figure 5: Acoustic Metamaterial (AMM) with a virtual ground. (a-i) Schematics of AMMs featuring a rigid inertial and a non-inertial ground. (a-ii) FFTs of velocity response, $\dot{v}_{\mathrm{d}}(\bar{t})$ in an AMM with rigid inertial grounding (red) and non-inertial grounding (blue). Both AMMs are subject to force $\mathrm{F}(\bar{t})$ (green), however, the rigid grounding case has no environmental noise while the non-inertial grounding case features a broadband environmental noise (orange) contaminating the AMM response via the grounding springs. (b-i) Schematic of a modified AMM where the last mass functions as a virtual ground, and isolates the AMM from undesirable external noise while simultaneously retaining the edge defect mode characteristics relatively intact. (b-ii) FFTs of $\dot{v}_{\mathrm{d}}(\bar{t})$ in an AMM with virtual grounding subject to force $\mathrm{F}(\bar{t})$. (c-i) Reduced order mass-spring model of the AMM with virtual grounding, where grounding springs connect all masses to the last mass (opposite boundary), (c-ii) eigenmode of the AMM at $\Omega_{\mathrm{d,AMM}}=0.8447$, and (c-iii) the convergence plot of the edge defect frequency, $\Omega_{\mathrm{d,AMM}}$.
As described in Sec. II, although grounding elements are not necessary for the existence of defect modes, their presence enables the existence of a defect mode as the principal eigenmode of a DMoA-based AMM within the bounded BG – $\Omega\in[0,\sqrt{k_{\mathrm{g,r}}})$. However, AMMs have been deployed in a variety of environments, e.g., rotating shafts[47], aerial vehicles[41], seismic structures[48] where a true inertial frame of reference may not be accessible. In this context, we present a modified edge DMoA-based AMM that can function unhindered without a rigid inertial reference by utilizing one of its boundaries as a virtual ground.
The schematics in Fig. 5(a,b-i) contrast AMM configurations with an inertial, non-inertial, and virtual ground. Fig. 5(a,b-ii) contrasts the FFTs of the edge defect velocities when subject to a Gaussian wave-packet type forcing, $\mathrm{F}(\bar{t})$ (centered at $\Omega_{\mathrm{d,DMoA}}$) in these three grounding configurations. The FFT (red) in the rigid grounding scenario reveals a behavior consistent with dynamics characterized by the edge DMoA model in Fig. 1b. However, the presence of broadband environmental noise (orange) renders the grounding mechanism non-inertial, causing the FFT (blue) in this scenario to deviate from its rigid grounding counterpart. Even though the environmental noise is chosen to have a maximum amplitude an order lower ($O(10^{-1})$) than the excitation signal, $\mathrm{F}(\bar{t})$, the impact on the AMM dynamics is evident.
Since the nature of the external grounding reference is dictated by the operating environment, we consider an alternate AMM configuration as shown in Fig. 5(b-i) to mitigate the impact of a non-inertial grounding reference. In the modified AMM, all constituent masses are now connected to the last mass which functions as an in-house virtual ground, isolating all AMM masses from the external environmental noise. Fig. 5(c-i) presents an equivalent mass-spring reduced order model for this altered AMM configuration with a virtual ground. An eigen-analysis (see Sec. A5) reveals completely new eigenvalues compared to the edge DMoA-based AMM, with three BG resonances – $\Omega=0$ and $\Omega_{\mathrm{d,AMM}}=0.8447$ in the low-frequency BG, and $\Omega=6.4095$ in the high-frequency BG, and the remainder lying within the PB range of the base MoA. Resonances at $\Omega=0$ and $\Omega_{\mathrm{d,AMM}}=0.8447$, respectively, correspond to the rigid body translation mode and the localized mode shape plotted in Fig. 5(c-ii) while the resonance at $\Omega=6.4095$ corresponds to a new eigenmode localized at the last mass, created as a result of the altered grounding spring connections. Though the eigenmode frequencies and shapes are significantly different from those of an edge DMoA-based AMM, the eigenmode at $\Omega_{\mathrm{d,AMM}}=0.8447$ (Fig. 5(c-ii)) emerges as an exception, existing relatively unaltered compared to the edge defect mode shown in the inset of Fig. 1b, owing to the last mass being a nodal point in the defect mode. The FFT curve (red) in Fig. 5(b-ii) captures the AMM dynamics in the presence of an in-house virtual ground which is identical to that of an AMM with a rigid ground (Fig. 5(a-ii)). Therefore, our modified analytical approach in conjunction with the presence of an in-house virtual ground provides an enhanced capability to design a DMoA-based AMM that can function efficiently at a prescribed BG resonance regardless of the nature of the operating environment.
A closer look at Fig. 5(b-ii) reveals a slight change in defect modal frequency, $\Omega_{\mathrm{d,AMM}}$ compared to the analytical DMoA edge defect frequency, $\Omega_{\mathrm{d,DMoA}}=0.8165$. This discrepancy stems from the limitation that the decaying eigenmode reaches a near-zero and not an absolute zero displacement at the last mass, i.e., $v_{\mathrm{N}}\approx 0$. Fig. 5(c-iii) illustrates this phenomenon, where choosing a larger AMM (large $\mathrm{N}$) aids in converging the displacement, $v_{\mathrm{N}}\rightarrow 0$ and the frequency, $\Omega_{\mathrm{d,AMM}}\rightarrow\Omega_{\mathrm{d,DMoA}}$.
III.2 Customized Acoustic absorber
Figure 6: Customized Acoustic absorber. (a) An edge DMoA-based AMM design with an engineered edge defect mode at $\Omega_{\mathrm{d}}=1$. (b) Velocity response, (c) the corresponding FFTs, and (d) Energy distribution at mass indices – (i) $j=1$, (ii) $j=(\mathrm{N}+1)/2$, and (iii) $j=\mathrm{N}$ ($\mathrm{N}=51$), of three different AMM configurations based on the base MoA (black), and two edge DMoAs with defect properties $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}\rightarrow$$\{1,0.382\}$ ($\square$), and $\{2.618,2\}$ ($\bullet$) used in Fig. 3.
Here we present an edge DMoA-based AMM that interacts with a broadband acoustic field, $\mathrm{F}(\bar{t})$, and manifests a prescribed narrow-band response at the edge defect while simultaneously customizing the transmission of PB frequencies through the AMM bulk. Fig. 6a shows the mass-spring configuration of an edge DMoA-based AMM subject to an external force, $\mathrm{F}(\bar{t})$ at the edge defect location. $\mathrm{F}(\bar{t})$ is chosen as a representative broadband acoustic signal (as illustrated by the green FFT curve in Fig. 6(c-i)) which can originate from a variety of environmental sources, e.g., vehicular traffic, fluid flow. A customized acoustic interaction of the AMM with this signal can be useful in mitigating or harvesting this ambient energy. To investigate this capability, we numerically simulate the response of the base MoA (black) devoid of material defects and two AMM configurations described by edge DMoA parameters chosen from Fig. 3 (blue and red). Fig. 6(b-d) plot, respectively, the velocity response of the three AMM configurations, the FFTs of these velocity signals (normalized by peak the largest FFT amplitude of all velocity signals), and the energy distribution at mass indices – (i) $j=1$ (edge defect), (ii) $j=(\mathrm{N}+1)/2$ (midpoint), and (iii) $j=\mathrm{N}$ (last mass).
The velocity responses and the corresponding FFTs of the base MoA in Fig. 6(b,c) indicate its behavior as a passive acoustic filter that allows propagation of frequency components of $\mathrm{F}(\bar{t})$ in the PB range, $\Omega\in[\sqrt{k_{\mathrm{g,r}}},\sqrt{4+k_{\mathrm{g,r}}}]$ while attenuating frequencies in the BGs. Alternately, both the AMMs manifest a narrow-band velocity response, centered at the prescribed defect frequency, $\Omega_{\mathrm{d}}=1$ at $j=1$. Though PB frequency signatures of $\mathrm{F}(\bar{t})$ persist at $j=(\mathrm{N}+1)/2,\mathrm{N}$ in the AMMs, the level of transmission is significantly lower compared to the base MoA. To analyze this contrasting behavior, we quantify the total energy at a given mass index:
$\mathrm{TE}_{j}=\mathrm{KE}_{j}+\mathrm{PE}_{j},$
where the kinetic and potential energy components are calculated as:
[…]
$\mathrm{PE}_{j}=\sum_{\bar{t}=0}^{\bar{t}=1000}\left\{\begin{array}[]{ll}\frac{1}{2}\left[k_{\mathrm{d}}(v_{\mathrm{d}}-v_{2})^{2}+k_{\mathrm{gd}}v_{\mathrm{d}}^{2}\right],&\mathrm{if}\ j=1,\\[8.5359pt] \hfill\frac{1}{2}\left[k(v_{j}-v_{j+1})^{2}+k_{\mathrm{g}}v_{j}^{2}\right],&\mathrm{if}\ j\in[2,\mathrm{N-1}],\\[8.5359pt] \hfill\frac{1}{2}\left[k_{\mathrm{g}}v_{\mathrm{N}}^{2}\right],&\mathrm{if}\ j=\mathrm{N},\end{array}\right.$
Consistent with the velocity responses in Fig. 6(b-ii,iii), the total energy bars, $\mathrm{TE}_{(\mathrm{N}+1)/2,\mathrm{N}}$ in Fig. 6(d-ii,iii), are significantly lower in both the AMMs compared to the base MoA indicating a reduction in transmission into the material bulk. This contrast in the transmission is facilitated by a superior energy localization at the edge defect in both AMMs, as evident from the higher energy content in Fig. 6(d-i). Furthermore, each AMM manifests a distinct energy localization behavior commensurate with the prescribed $\dot{\mathrm{V}}_{\mathrm{E}}$ scales from Fig. 3. The persistence of these prescribed velocity scales throughout the AMMs (Fig.6(b,d-ii,iii)) reiterates the importance of the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ as a critical behavioral parameter that customizes both the localized response at the defect location and the energy transmission into the bulk.
Additional modifications to the AMM, e.g., adding more defects[26], secondary branching of the mass-spring chain[49], can be made according to operational requirements to further customize the AMM behavior without affecting the characteristics of the existing edge defect under reasonable proximity limits(II.3). Therefore, our modified analytical method offers a systematic strategy to design multi-functional defect-embedded AMMs based on simple mass-spring reduced-order models.
IV Conclusion
In summary, we present a modified analytical approach that allows us to engineer BG resonances with prescribed characteristics in AMMs by embedding material defects and breaking the periodicity of the underlying material matrix. The perturbed Toeplitz matrix eigenvalue problem informs our analytical approach which accurately quantifies the acoustic characteristics of single and multiple defect-embedded AMMs with reasonable proximity limits. Furthermore, we demonstrate the benefits of defect modes through two DMoA-based AMM examples that feature an in-house virtual ground, and a customized acoustic absorption of an incident broadband acoustic field. The proposed strategy can be readily extended to lattices operating under general boundary conditions[28, 30] and polyatomic unit-cell[28, 29] based lattices as required, and augments the toolkit for designing defect-based AMMs.
CRediT authorship contribution statement
Vinod Ramakrishnan (VR): conceptualization, formal analysis, methodology, writing – original draft, review, and editing. Kathryn Matlack (KM): supervision, formal analysis, funding acquisition, writing – review, and editing.
Data availability
Data will be made available on request.
Acknowledgments
VR and KM acknowledge funding from AFOSR MURI under grant number FA9550-23-1-0299.
Appendix
A1 Edge defect eigenmode
The matrix equation of motion for an edge defect-embedded grounded monoatomic chain (edge DMoA), with the temporal frequency normalized by $\omega_{0}=\sqrt{k/m}$, is given as:
$\mathbf{M}\ddot{\mathbf{V}}+\mathbf{K}\mathbf{V}=\mathbf{0}$
where,
$\mathbf{M}=\begin{bmatrix}m_{\mathrm{r}}&0&\dots&0\\ 0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&1\end{bmatrix}_{\mathrm{N}\times\mathrm{N}},\mathbf{K}=\begin{bmatrix}k_{\mathrm{r}}+k_{\mathrm{gd,r}}&-k_{\mathrm{r}}&0&0&\dots&0\\ -k_{\mathrm{r}}&1+k_{\mathrm{r}}+k_{\mathrm{g,r}}&-1&0&\dots&0\\ 0&-1&2+k_{\mathrm{g,r}}&-1&\dots&0\\ 0&0&-1&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&\ddots&&-1\\ 0&0&0&\dots&-1&1+k_{\mathrm{g,r}}\end{bmatrix}_{\mathrm{N}\times\mathrm{N}},$ (A1)
and, $m_{\mathrm{r}}=m_{\mathrm{d}}/m$, $k_{\mathrm{r}}=k_{\mathrm{d}}/k$, $k_{\mathrm{g,r}}=k_{\mathrm{g}}/k$ and $k_{\mathrm{gd,r}}=k_{\mathrm{gd}}/k$ represent the defect mass ratio, defect interaction stiffness ratio, grounding spring stiffness ratio and the defect grounding spring stiffness ratio, respectively. Note that defect eigenmodes exist for cases where, $k_{\mathrm{r}}\neq 1$ however, the procedure to analytically estimate the eigenfrequencies is challenging when introducing a discrepancy in off-diagonal terms of the stiffness matrix. Hence, we set $k_{r}=1$ for all cases explored in this article to obtain a stiffness matrix conducive to formulating analytical solutions for the matrix eigenvalue problem.
Assuming harmonic motion, $\mathbf{V}=\hat{\mathbf{V}}e^{i\Omega\bar{t}}$, the eigenvalue problem for the DMoA can be formulated as $|\mathbf{K}_{\mathrm{d}}|=\mathbf{0}$, where,
$\mathbf{K}_{\mathrm{d}}=\mathbf{K}-\Omega^{2}\mathbf{M}=\begin{bmatrix}d+\alpha&a&0&0&\dots&0\\ a&d&a&0&\dots&0\\ 0&a&d&a&\dots&0\\ 0&0&a&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&\ddots&&a\\ 0&0&0&\dots&a&d+\beta\end{bmatrix}$ (A2)
is the dynamic stiffness matrix and the matrix coefficients are given as,
$d=2+k_{\mathrm{g,r}}-\Omega^{2},a=-1,\beta=-1,$ (A3)
$\alpha=(k_{\mathrm{gd,r}}-k_{\mathrm{g,r}})-1+(1-m_{\mathrm{r}})\Omega^{2};$
The tridiagonal matrix, $\mathbf{K}_{\mathrm{d}}$ is consistent with the general form of a perturbed tridiagonal Toeplitz matrix[43, 28]. Without loss of generality, assuming an odd number of degrees of freedom in the system, i.e., $\mathrm{N}=2n+1$, the characteristic equation corresponding to an eigenvalue problem of a perturbed tridiagonal Toeplitz matrix is given by:
$(-d-\alpha-\beta)P_{n}-[\alpha\beta d+(\alpha+\beta)a^{2}]P_{n-1}=0$ (A4)
where,
$P_{n}=P_{n}(x)=a^{2n}U_{n}(x)$
and, $U_{n}(x)$ is the Chebyshev polynomial of the second kind,
$U_{n}(x)=\frac{\sin(2(n+1)\kappa)}{\sin(2\kappa)},$ (A5)
$x=\cos(2\kappa)=\frac{d^{2}}{2a^{2}}-1$
In the above equations, $\kappa$, represents the spatial frequency (non-dimensional wavenumber) associated with a given harmonic temporal frequency, $\Omega$. Eq. (A4) is an implicit equation that cannot be solved to obtain an explicit analytical form relating the DMoA material parameters to $\kappa$ and $\Omega$. Hence, we rearrange the LHS in Eq. (A4) to generate the modified equation, Eq. (3) listed in the main text.
Figure A1: Graphical representation of functions $f(\Omega)$ and $g(\Omega)$ from Eq. (3) in the main text.
Eq. (3) is solved numerically by sweeping through a given range of frequencies, $\Omega$ to identify points of intersections between functions $f$ and $g$. A graphical representation of the same is shown in Fig. A1. Despite the U-transformation method utilized by Cai et al.[32] for the case of a bulk defect and the perturbed Toeplitz matrix-based method that we propose for the edge defect being fundamentally different solution approaches, a noteworthy similarity between the procedures is that they both involve implicit equations in the final step that need to be solved numerically to obtain the defect eigenfrequencies and eigenmode shapes.
A2 Multiple BG resonances with single edge defect
Though, the design space of the defect properties, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ probed in Fig. 2 in the main text manifest a single defect mode, other combinations, especially cases where $m_{\mathrm{r}}<1$ and $k_{\mathrm{gd,r}}<k_{\mathrm{g,r}}$ can manifest an additional defect mode in $2^{\mathrm{nd}}$ BG of the MoA, localized at the same edge defect location as the existing defect mode. Fig. A2 presents the (a) defect frequencies, $\Omega_{\mathrm{d1,d2}}$ and (b) the decay rates, $\kappa_{I}$ when probing the defect material properties, $m_{\mathrm{r}}\in[0.04,1]$ and $k_{\mathrm{gd,r}}=0$. Not only do the analytical and numerical values of the defect frequencies and decay rates agree well, but we also notice that below the value, $m_{\mathrm{r}}\approx 0.33$ the eigenfrequencies of the finite DMoA undergo a bifurcation, where the highest PB frequency relocates into the $2\mathrm{nd}$ BG, appearing as a second defect resonance frequency in the DMoA chain.
Figure A2: Transition of a DMoA with material defect at a single boundary from manifesting a single defect mode to two defect modes. (a) Defect frequencies, $\Omega_{\mathrm{d1,d2}}$ and, the corresponding (b) decay rates, $\kappa_{I}$. ($k_{\mathrm{gd,r}}=0$)
A3 Resonance Velocity Amplitude Envelope
The matrix equation governing the forced response of a DMoA subject to an external, single-frequency harmonic forcing is given as:
$\mathbf{M}\ddot{\mathbf{V}}+\mathbf{K}\mathbf{V}=\frac{\mathbf{F}_{0}}{2}e^{i\Omega_{\mathrm{exc}}\bar{t}}+\mathrm{c.c.},$ (A6)
where, $\mathbf{F}_{0}$ represents an $\mathrm{N}\times 1$ column vector with complex force excitation amplitudes (i.e., $\mathbf{F}_{0,j}\in\mathbb{C}$ for $j=1,2,\cdots,\mathrm{N}$), $\Omega_{\mathrm{exc}}$ represents the non-dimensional excitation frequency, and $\mathrm{c.c.}$ stands for the complex conjugate terms. If the excitation frequency coincides with one of the natural frequencies of the systems, i.e., $\Omega_{\mathrm{exc}}=\Omega_{j}$, the general solution ansatz for the displacement of the DMoA at resonance is of the form:
$\mathbf{V}(\bar{t})=\hat{\mathbf{V}}_{j}\left(\frac{A+B\bar{t}}{2}\right)e^{i\Omega_{j}\bar{t}}+\mathrm{c.c.},$ (A7)
where, $\hat{\mathbf{V}}_{j}$ represents the normalized eigenvector (i.e., $\hat{\mathbf{V}}_{j}^{\mathrm{T}}\hat{\mathbf{V}}_{j}=1$) corresponding to $\Omega_{j}$ and coefficients $A,B\in\mathbb{C}$. Specifically, for $\Omega_{j}=\Omega_{\mathrm{d}}$ and, initial displacement and velocity conditions - $\mathbf{V}(0)=\mathbf{0}$ and $\dot{\mathbf{V}}(0)=\mathbf{0}$, the eigenvector, $\hat{\mathbf{V}}_{\mathrm{d}}$ and the coefficients $A,B$ can be written as:
$\hat{\mathbf{V}}_{\mathrm{d}}=\widetilde{\mathrm{V}}_{\mathrm{d}}\mathbf{V}_{\mathrm{d}}=\widetilde{\mathrm{V}}_{\mathrm{d}}e^{i\kappa_{R}\mathbf{x}}e^{-\kappa_{I}\mathbf{x}},$ (A8a)
$\mathbf{V}(0)=\mathbf{0}\Longrightarrow$ $\mathrm{Re}(A)=0$
$\Longrightarrow$ $A=ia,\mathrm{for}\ a\in\mathbb{R},$ (A8b)
$\dot{\mathbf{V}}(0)=\mathbf{0}\Longrightarrow$ $\mathrm{Re}(B)=\Omega_{\mathrm{d}}\mathrm{Im}(A)$
$\Longrightarrow$ $B=a\Omega_{\mathrm{d}}+ib,\mathrm{for}\ a,b\in\mathbb{R},$ (A8c)
where, $\mathbf{x}=[0,1,\cdots,\mathrm{N}-1]^{\mathrm{T}}$ represents the coordinate vector, $\kappa_{R}$ and $\kappa_{I}$, respectively, represent the real and imaginary wavenumber (decay rate), and $\widetilde{\mathrm{V}}_{\mathrm{d}}$ represents the normalizing factor associated with the defect eigenmode, $\hat{\mathbf{V}}_{\mathrm{d}}$. Assuming a defect mode in the $1^{\mathrm{st}}$ BG ($2^{\mathrm{nd}}$ BG), i.e., $\kappa_{R}=0$ ($\kappa_{R}=\pi$) and, substituting, Eq. (A7) and Eq. (A8) in Eq. (A6), we get:
$\left\{\left(\frac{\Omega_{\mathrm{d}}}{2}\right)\left[a\bar{t}(\mathbf{K}-\Omega_{\mathrm{d}}^{2}\mathbf{M})-2b\mathbf{M}\right]+\left(\frac{i}{2}\right)\left[\mathbf{K}(a+b\bar{t})+\mathbf{M}(a-b\bar{t})\Omega_{\mathrm{d}}^{2}\right]\right\}\hat{\mathbf{V}}_{\mathrm{d}}e^{i\Omega_{\mathrm{d}}\bar{t}}+\mathrm{c.c.}=\frac{\mathbf{F}_{0}}{2}e^{i\Omega_{\mathrm{d}}\bar{t}}+\mathrm{c.c.},$ (A9)
The real and imaginary parts of the coefficients of $e^{i\Omega_{\mathrm{d}}\bar{t}}$ on the LHS and RHS in Eq. (A9) can be equated to obtain two matrix equations:
$\left[a\bar{t}(\mathbf{K}-\Omega_{\mathrm{d}}^{2}\mathbf{M})-2b\mathbf{M}\right]\hat{\mathbf{V}}_{\mathrm{d}}=\frac{\mathrm{Re}(\mathbf{F}_{0})}{\Omega_{\mathrm{d}}};$ (A10a)
$\left[\mathbf{K}(a+b\bar{t})+\mathbf{M}(a-b\bar{t})\Omega_{\mathrm{d}}^{2}\right]\hat{\mathbf{V}}_{\mathrm{d}}=\mathrm{Im}(\mathbf{F}_{0}),$ (A10b)
Now, assume an external sinusoidal forcing, $\mathbf{F}=[\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t}),0,\cdots,0]^{\mathrm{T}}$, applied on the DMoA boundary housing the (edge) defect. The complex forcing amplitudes are expressed as:
$\mathrm{Re}(\mathbf{F}_{0})=\mathbf{0},\mathrm{Im}(\mathbf{F}_{0})=[-\mathrm{F}_{0},0,\cdots,0]^{\mathrm{T}},$
Substituting, the above equation into Eq. (A10b)a, we get $b=0$, as $\Omega_{\mathrm{d}}$ and $\hat{\mathbf{V}}_{\mathrm{d}}$ are a solution to the eigenvalue problem $\mathbf{K}_{\mathrm{d}}\hat{\mathbf{V}}_{\mathrm{d}}=0$. Therefore, the DMoA response to the given sinusoidal forcing can be written as:
$\mathbf{V}$ $=\left(\frac{\dot{\mathrm{V}}_{\mathrm{E}}}{\Omega_{\mathrm{d}}^{2}}\right)\left[\sin(\Omega_{\mathrm{d}}\bar{t})-\Omega_{\mathrm{d}}\bar{t}\cos(\Omega_{\mathrm{d}}\bar{t})\right]\mathbf{V}_{\mathrm{d}},$ (A11)
$\dot{\mathbf{V}}$ $=\left[\dot{\mathrm{V}}_{\mathrm{E}}\bar{t}\sin(\Omega_{\mathrm{d}}\bar{t})\right]\mathbf{V}_{\mathrm{d}},$
where, Eq. (A10b)b yields:
$\dot{\mathrm{V}}_{\mathrm{E}}=-\Omega_{\mathrm{d}}^{2}\widetilde{\mathrm{V}}_{\mathrm{d}}\left[\hat{\mathbf{V}}_{\mathrm{d}}^{\mathrm{T}}\left[\mathbf{K}+\mathbf{M}\Omega_{\mathrm{d}}^{2}\right]^{-1}\mathrm{Im}(\mathbf{F}_{0})\right],$ (A12)
A4 Multiple defects
Figure A3: Multi-DMoA chain with defect frequencies in $2^{\mathrm{nd}}$ BG of the MoA. (a) Absolute percentage error contours for a Multi-DMoA ($\mathrm{N}=21$) with two defects at $x_{\mathrm{d1}},x_{\mathrm{d2}}\in[0,\mathrm{N}-1]$ and $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$, (b) Eigenmode shapes of the Multi-DMoA with one edge ($x_{\mathrm{d1}}=0$) and one bulk ($x_{\mathrm{d2}}$) defect. Multi-DMoA material parameters - $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$, $k_{\mathrm{r}}=1$ and $\{m_{r},k_{\mathrm{gd,r}}\}=\{1,6\}$
The modified analytical approach presented in this article can be applied to determine the defect resonance frequencies of a Multi-DMoA provided the additional defects remain relatively dormant at the resonance frequency of a considered defect. To verify the accuracy of the analytical estimate of $\Omega_{\mathrm{d}}$, Fig. 4 in the main text plots the absolute percentage error b.w. the numerical and analytical values of the defect frequency given by,
%$\%\ \mathrm{Error}\ \mathrm{in}\ \Omega_{\mathrm{d}}=\frac{\Delta\Omega_{\mathrm{d}}}{\Omega_{\mathrm{d},\mathrm{An.}}}\times 100\%,$ (A13)
where, $\Delta\Omega_{\mathrm{d}}=\left|\Omega_{\mathrm{d},\mathrm{Num.}}-\Omega_{\mathrm{d},\mathrm{An.}}\right|$.
When the defect frequencies lie in the $1^{\mathrm{st}}$ BG of the MoA, we notice a small surge in the $\Omega_{\mathrm{d1}}$ error when at least one of the penultimate masses are prescribed defects, i.e, $x_{\mathrm{d1,d2}}=1\ \mathrm{or}\ \mathrm{N}-2$. This error surge is attributed to the spatial in-phase vibration associated with the defect modes in the $1^{\mathrm{st}}$ BG and the significant impact of defect proximity to the finite boundary. Alternately, Fig. A3 explores a Multi-DMoA chain with defect modes now in the $2^{\mathrm{nd}}$ BG of a MoA chain. The error contours at $x_{\mathrm{d1,d2}}=1\ \mathrm{or}\ \mathrm{N}-2$ in Fig. A3a are virtually indistinguishable from the surrounding configurations (blue region - $0.003\%$, surge region - $0.29\%$) as opposed to the error contours in Fig. 4a (blue region - $0.004\%$, surge region - $3.9\%$). The reduction in the error surge is indeed a result of improved accuracy of the analytical model and out-of-phase vibration of the defect and boundary masses as the absolute error, $\Delta\Omega_{\mathrm{d}}$ also decreases for defect modes in the $2^{\mathrm{nd}}$ BG.
A5 AMM with a Virtual Ground
The modified stiffness matrix,
$\mathbf{K}_{\mathrm{VG}}=\begin{bmatrix}k_{\mathrm{r}}+k_{\mathrm{gd,r}}&-k_{\mathrm{r}}&0&0&\dots&-k_{\mathrm{gd,r}}\\ -k_{\mathrm{r}}&1+k_{\mathrm{r}}+k_{\mathrm{g,r}}&-1&0&\dots&-k_{\mathrm{g,r}}\\ 0&-1&2+k_{\mathrm{g,r}}&-1&\dots&-k_{\mathrm{g,r}}\\ 0&0&-1&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&\ddots&&-1-k_{\mathrm{g,r}}\\ -k_{\mathrm{gd,r}}&-k_{\mathrm{g,r}}&-k_{\mathrm{g,r}}&\dots&-1-k_{\mathrm{g,r}}&1+(\mathrm{N}-2)k_{\mathrm{g,r}}+k_{\mathrm{gd,r}}\end{bmatrix}_{\mathrm{N}\times\mathrm{N}},$
and $\mathbf{M}=diag([m_{\mathrm{r}},1,1,\cdots,1]_{\mathrm{N}\times 1})$ can be substituted in Eq. (A2) to calculate the finite eigenvalues of the AMM with a virtual ground.